Solve for x:
4x – 7 = -3(6 – 1)

Answer by checkley71(8403)

4X-7=-3(6-1) OR 4X-7=-3*5 OR 4X-7=-15 OR 4X=-15+7 OR 4X=-8 OR X=-8/4 OR X=-2
PROOF 4(-2)-7=-3(6-1) OR -8-7=-3(5) OR -15=-15